Write the augmented matrix for the system of equations and solve the system using Gauss-Jordan Elimination.
10x+8y+20Z=6,000
10x+20y+40Z=11,000
100X+20y+200z=36,900
Help PLEASE :)

Answer by stanbon(48568)

Write the augmented matrix for the system of equations and solve the system using Gauss-Jordan Elimination.
----------------------------------
10x+8y+20Z=6,000
10x+20y+40Z=11,000
100X+20y+200z=36,900
--------------------
Subtract 1st from 2nd and Subtract 10 times 1st from 3rd to get:
10x+8y+20Z=6,000
0x+12y+20z = 5000
0x-60y+0z = -23100
---------------------
Solve the 3rd equation for "y"
y = 385
-------------
Substitute into the 2nd and solve for "z"
12*385 + 20z = 5000
z = 19
-------------------
Substitute into the 1st and solve for "x"
10x + 8*385 + 20*19 = 6000
x = 254
=================
Cheers,
Stan H.